# WW-FL: Secure and Private Large-Scale Federated Learning

### v3

@@ Abstract.
Federated Learning, Secure Multi-party Computation, Privacy, MPC, Hierarchical FL, Poisoning, Attacks and Defenses
1. Introduction
@@
Federated learning (FL) (McMahan et al., 2017) is a revolutionary approach for large-scale distributed machine learning that emerged in the past decade and has had a profound impact on both academia (Yin et al., 2022) and industry (Rieke et al., 2020; Ludwig et al., 2020; He et al., 2020). A World Wide Web Consortium (W3C) community group currently works towards establishing FL-related Web standards (W3C, 2023). FL shares some objectives with privacy-preserving machine learning (PPML) (Gilad-Bachrach et al., 2016; Mohassel and Zhang, 2017; Cabrero-Holgueras and Pastrana, 2021), but differs in its approach to training a model: In PPML, training on the entire set of available data is either outsourced to a remote party or done collaboratively while ensuring data privacy. In contrast, FL involves multiple rounds of training, where each round consists of selected clients locally training a model on their private data and a central party performing aggregation of these models to update a global model.
@@
Initially, FL appeared to indeed offer privacy since the training data remains on the client devices, and only model updates (so-called gradients) are transmitted. However, subsequent research has demonstrated that these gradients still reveal a significant amount of information, enabling the central aggregator to infer sensitive details about the clients’ private training data (Zhu et al., 2019; Geiping et al., 2020). To address this issue, secure aggregation (SA) schemes (Bonawitz et al., 2017; Bell et al., 2020) were proposed, where only the aggregated result is revealed to the central party, effectively concealing the individual gradients. Unfortunately, FL poses additional challenges that require careful attention.
Figure 1. Visual representation of deploying our secure world-wide FL (WW-FL) framework. In this illustration, clients connect to local MPC clusters, which in turn are connected to three global servers (this can be varied as per requirements) strategically positioned across continents. Fig. 2 provides a detailed illustration of WW-FL, showcasing the roles of involved entities.
@@
The first challenge relates to serious privacy vulnerabilities that have been identified when using secure aggregation with a single aggregator (So et al., 2021; Boenisch et al., 2021, 2022; Fowl et al., 2022; Wen et al., 2022): Recent studies like Mandrake (Zhao et al., 2023) have demonstrated the effectiveness of privacy attacks even when utilizing secure aggregation techniques across a large number of clients, showing that such schemes are insufficient (Pasquini et al., 2022; Boenisch et al., 2023). This hinders utilizing standard FL and even FL with SA for global, worldwide training tasks as sharing sensitive data across jurisdictions in many cases is legally problematic. For example, the Court of Justice of the European Union (CJEU) with its famous “Schrems II” judgment has ruled the “Privacy Shield” agreement between the EU and the US for exchanging personal data invalid (Tracol, 2020). Until agreeing on the most recent “Data Privacy Framework” (U.S. Department of Commerce, 2022), numerous attempts to set up data sharing have failed, including the “Safe Harbour” principles (Weiss and Archick, 2016). Recent works therefore suggest a distributed aggregator setup based on secure multi-party computation (MPC) techniques to securely distribute aggregation tasks among multiple servers, ensuring privacy despite potential collusion among a subset of them (Fereidooni et al., 2021; Ben-Itzhak et al., 2022; Rathee et al., 2023; Gehlhar et al., 2023).
@@
The second (and closely related) challenge is the current lack of global model privacy. Initially, FL focused on enhancing user participation and improving model accuracy rather than safeguarding the model itself. However, with the widespread adoption of FL in various industries, such as the medical sector (Rieke et al., 2020), the need for privacy protection becomes apparent. Unfortunately, unrestricted access to the aggregated model allows extraction of traces of original training data, necessitating global model privacy (Pasquini et al., 2022; Boenisch et al., 2023; Zhao et al., 2023). So far, only a small number of works have focused on preserving the privacy of the global model in FL (Mandal and Gong, 2019; Sav et al., 2021). Most of these works use homomorphic encryption (HE) schemes (Acar et al., 2018), which primarily address semi-honest corruptions. Moreover, several works do not consider the collusion of clients with the aggregator and thereby provide a weaker notion of security.
@@
The third challenge arises when corrupted clients intentionally manipulate the locally trained model to reduce the system’s accuracy or to introduce backdoors to extract sensitive information in future rounds (Wang et al., 2020; Xie et al., 2020). This method of model poisoning is more potent than data poisoning (Shejwalkar et al., 2022; Xie et al., 2024), which involves manipulating the underlying dataset. Although numerous defenses have been proposed against model poisoning attacks (Shejwalkar and Houmansadr, 2021; Cao et al., 2021), the most effective ones remain impractical for large-scale deployment due to their significant computation and communication costs (Nguyen et al., 2022). Additionally, recent research has introduced even more potent model poisoning strategies that can circumvent state-of-the-art defenses (Shejwalkar and Houmansadr, 2021; Xie et al., 2024). For instance, PoisonedFL (Xie et al., 2024) breaks eight state-of-the-art defenses without any knowledge of the honest clients’ models, highlighting the urgent need for new defense mechanisms. To address corrupted clients, we limit their attack capabilities to data poisoning rather than the more severe model poisoning. This allows us to implement more moderate defenses compared to existing schemes (cf. §5).
@@
Besides these three challenges, a large-scale, worldwide FL deployment needs to account for client heterogeneity in terms of bandwidth, jurisdiction, trust assumptions, and computational power. To elaborate, trust in government and private entities varies widely, influencing how data is shared and used. For instance, strict data privacy laws in some countries impose heavy fines for non-compliance, while data breaches in companies like TalkTalk [^1] https://www.talktalkgroup.com and Target [^2] https://target.com/ have led to significant losses in customer trust and revenue (Erin, 2022; Hibberd, 2022). In the context of training machine learning models for sensitive applications, such as medical diagnostics, diverse patient data is crucial but must be handled with strict privacy.
@@
Although various works have addressed the aforementioned issues individually or in combination, a comprehensive solution remains elusive. Therefore, we propose a unified framework called WW-FL, which enables secure and private distributed machine learning at scale. Departing from traditional FL, we introduce the concept of a trust zone, allowing data to be shared securely within a trusted legal jurisdiction and leveraging privacy-enhancing technologies like MPC. This approach supports edge nodes with limited resources and encourages wider participation, enhancing the system’s overall privacy and performance. We give a more detailed discussion and additional motivation in App. A.1.
Our Contributions
@@
In this paper, we initiate a discussion on novel approaches towards secure and private large-scale federated learning to address the challenges outlined above. Towards this goal, we propose a unified framework called “worldwide-FL” (WW-FL). Our framework is based on a novel abstraction that also captures existing hierarchical FL architectures in a hybrid manner as visualized in Fig. 1 and detailed in §3.5.
Table 1. Comparison of WW-FL and previous works. Notations: $\mathcal{S}$ – Aggregation Server(s), $\mathcal{C}$ – Client, GM – Global Model, LM – Local Model, MPC – Secure Multi-party Computation, DP – Differential Privacy. Given the extensive literature, this comparison focuses on one representative work per category.
Categories Representative Work Method Privacy ($\mathcal{S}$) Privacy ($\mathcal{C}$) Defense Cross Device No Client Interaction Dropout Handling
GM LM GM
Aggregation (Plain) (McMahan et al., 2017)
Aggregation (Robust) (Yin et al., 2018)
Secure Aggregation (Single $\mathcal{S}$) (Bonawitz et al., 2017) Masking
Secure Aggregation (Multi $\mathcal{S}$) (Nguyen et al., 2022) MPC
Hierarchical FL (HFL) (Yang, 2021) DP
Worldwide FL (WW-FL) This work MPC
@@
Briefly, in our framework, FL clients use secret-sharing to securely outsource training data to distributed cluster servers that are based on MPC. The clients then can leave and only sporadically return to provide more training data – this makes our framework robust against real-world issues such as drop-outs, requires no interaction between clients, and resource-constrained (mobile or edge) devices have very little workload. The models, which are trained by the cluster servers with MPC-based PPML techniques, are then aggregated across all training clusters using one or multiple levels of distributed aggregators. For secure distributed aggregation, we again utilize MPC. Note that after aggregation, models are not publicly released but in secret-shared form handed back to the training clusters for the next training iteration. After training is completed, known secure inference protocols can be used to allow private queries (Mann et al., 2023) in a controlled (potentially rate-limited) way. This architecture design addresses issues with single aggregators by relying on a distributed aggregator and notably achieves global model privacy due to a novel combination of FL with PPML techniques.
@@
We observe that a unique property of our framework is the strictly limited attack surface: similar to PPML, malicious clients are restricted to weaker data-poisoning attacks as there is no possibility to access and manipulate the model itself. As defending against data poisoning in PPML is highly non-trivial, we show that state-of-the-art data-poisoning attacks in the suggested hierarchical configuration are less effective than in plain FL. Furthermore, we consider different robust aggregation schemes to further mitigate the effect of such attacks. For this, we additionally propose new heuristics that improve the efficiency in MPC. Specifically, we propose a lightweight variant of the “Trimmed Mean” (Yin et al., 2018) defense for our setting with only data-poisoning attacks.
@@
Finally, we systematically evaluate possible instantiations of WW-FL in terms of performance (both in terms of achievable accuracy as well as overhead due to MPC) and resilience against poisoning attacks. For this, we implement a prototype based on PyTorch and Meta’s CrypTen framework (Knott et al., 2021) for instantiating MPC. We then evaluate the performance when training neural networks for standard image classification tasks in realistic network settings and using GPU-accelerated AWS EC2 instances. This way, we show that the computation and communication overhead induced by MPC is feasible even for large-scale deployments. For example, one round of training LeNet privately with security against semi-honest servers on the MNIST dataset for 5 epochs with batch size 80 takes less than 5 minutes for a cluster with 10 clients, and global aggregation among 10 clusters takes less than 1 second. Moreover, we find that the impact of MPC-induced inaccuracies is less than 0.1%, and models trained with WW-FL converge significantly faster than in plain FL.
@@
    We propose a new generic (hierarchical) FL framework called WW-FL that achieves complete model privacy, supports resource-constrained mobile or edge devices, and significantly limits the attack surface for malicious clients to data poisoning only.
@@
In Tab. 1, we summarize how WW-FL distinguishes itself from related works.
2. Preliminaries and Related Work
@@
This section provides a brief background on privacy-enhancing technologies such as MPC and HE, as well as an introduction to FL, along with a concise overview of related works.
@@
Secure Multi-Party Computation (MPC). MPC (Yao, 1986; Goldreich et al., 1987) enables a set of mutually distrusting parties to evaluate a public function $f()$ on their private data while preserving input data privacy. The corruption among the parties is often modelled via an adversary that may try to take control of the corrupted parties and coordinate their actions. There are various orthogonal aspects of adversarial corruption like honest vs. dishonest majority, semi-honest vs. malicious corruption, etc. (Evans et al., 2018; Lindell, 2020). MPC is practically efficient when run among a small number of computing parties (Patra and Suresh, 2020; Knott et al., 2021; Keller and Sun, 2022; Koti et al., 2022). In our prototype implementation of WW-FL, we utilize Meta’s CrypTen framework (Knott et al., 2021), which for efficiency reasons operates in a semi-honest two-party setting with a trusted third “helper” party that generates correlated randomness (Riazi et al., 2018; Chaudhari et al., 2019; Patra et al., 2021).
@@
Homomorphic Encryption (HE). HE schemes (Rivest et al., 1978; Gentry, 2009) enable computation on encrypted data without the need for decryption. Additively homomorphic encryption (AHE) is a widely used method that allows for the generation of a new ciphertext representing the sum of multiple plaintexts through operations on their corresponding ciphertexts (Paillier, 1999). In scenarios involving multiple parties, recent multi-party homomorphic encryption (MHE) schemes can reduce the communication complexity of MPC, but have high computation overhead (Mouchet et al., 2021). Furthermore, they are often secure only against semi-honest corruptions, whereas the recent work in (Chatel et al., 2023) is an exception. A comprehensive survey of various HE schemes is given in (Acar et al., 2018).
@@
Differential Privacy (DP). The concept of DP (Dwork and Roth, 2014) is based on the idea of adding noise to data in order to reduce information leakage when sharing it, while still allowing for meaningful computations to be carried out on the data. Though DP techniques offer some protection against attacks in federated learning, they also reduce the utility of the data and the resulting ML model. Additionally, achieving robust and accurate models necessitates significant privacy budgets, leaving the actual level of privacy achieved in practice uncertain (Ouadrhiri and Abdelhadi, 2022). We refer to (Desfontaines and Pejó, 2020) for more details on DP.
@@
Federated Learning (FL). Unlike conventional PPML techniques, FL (Konečnỳ et al., 2016; McMahan et al., 2017) enables the training of a machine learning model on distributed data by allowing each device to train the model locally using its own data. In each iteration, these local models are sent to a central server (also called an aggregator), where they are aggregated to form a global model. At a high level, an FL scheme iterates through the following steps:
@@
Secure Aggregation (SA). Standard FL’s privacy issues (Zhu et al., 2019; Geiping et al., 2020) led to single-server secure aggregation protocols (Single $\mathcal{S}$(Bonawitz et al., 2017; Mansouri et al., 2023) that use masking techniques or HE (Rubin, 1996; Bonawitz et al., 2017; Bell et al., 2020; Phong et al., 2018; Mandal and Gong, 2019). However, complete privacy for the global model remains elusive, posing risks if clients and aggregators collude. Moreover, recent research has shown that a single malicious server can reconstruct individual training data points from users’ local models even when using secure aggregation protocols (Pasquini et al., 2022; Boenisch et al., 2023; Zhao et al., 2023). Such vulnerabilities against colluding parties are addressed by multi-server SA (Multi $\mathcal{S}$(Fereidooni et al., 2021; Dong et al., 2021; Nguyen et al., 2022) and techniques like Spindle and Poseidon (Froelicher et al., 2021; Sav et al., 2021), albeit with scalability challenges. Recent works like (Dong et al., 2021) and (Nguyen et al., 2022) combine HE and MPC techniques to achieve private and robust FL, albeit at the cost of considerable computation and communication overhead in cross-device scenarios comprising hundreds to thousands of clients (Shejwalkar et al., 2022). In WW-FL, we employ a multi-server aggregation scheme, and mitigate scalability issues of existing works, especially for a heterogeneous cross-device setting.

### v4

@@ Abstract
keywords:
Federated Learning Hierarchical FL MPC Poisoning
1 Introduction
@@
Federated learning (FL) [MMR+17] is a revolutionary approach for large-scale distributed machine learning that emerged in the past decade and has had a profound impact on both academia [YZH22] and industry [RHL+20, LBT+20, HLS+20]. A World Wide Web Consortium (W3C) community group currently works towards establishing FL-related Web standards [W3C25]. FL shares some objectives with privacy-preserving machine learning (PPML) [GDL+16, MZ17, NC23, SBBE25], but differs in its approach to training a model: In PPML, training on the entire set of available data is either outsourced to a remote party or done collaboratively while ensuring data privacy. In contrast, FL involves multiple rounds of training, where each round consists of selected clients locally training a model on their private data and a central party performing aggregation of these models to update a global model.
@@
Initially, FL appeared to indeed offer privacy since the training data remains on the client devices, and only model updates (so-called gradients) are transmitted. However, subsequent research has demonstrated that these gradients still reveal a significant amount of information, enabling the central aggregator to infer sensitive details about the clients’ private training data [GBDM20, ZSE+24, WBS+24, DHW+25]. To address this issue, secure aggregation (SA) schemes [BIK+17, BBG+20, MOJC23] were proposed, where only the aggregated result is revealed to the central party, effectively concealing the individual gradients. Despite this, FL still poses several complex challenges that demand careful consideration [ZBA+25].
The first challenge relates to serious privacy vulnerabilities that have been identified when using secure aggregation with a single aggregator [BDS+22, FGC+22, WGF+22, SAG+23, BDS+23b, ZSE+24]: Recent studies have demonstrated the effectiveness of privacy attacks even when utilizing secure aggregation techniques across many clients, showing that such schemes are insufficient [PFA22, BDS+23a]. These vulnerabilities largely stem from the imbalance of power held by the central aggregator, who can introduce inconsistencies in the global model, manipulate client selection, and exert disproportionate influence over the training process.
@@
These growing concerns over privacy with a single central aggregator hinder the use of standard FL, and even FL with SA, for global training tasks as sharing sensitive data across jurisdictions is often problematic legally. For example, the Court of Justice of the European Union (CJEU) with its famous “Schrems II” judgment has ruled the “Privacy Shield” agreement between the EU and the US for exchanging personal data invalid [Tra20]. Until agreeing on the most recent “Data Privacy Framework” [U.S22], numerous attempts to set up data sharing have failed, including the “Safe Harbour” principles [WA16]. Recent works therefore suggest a distributed aggregator setup using secure multi-party computation (MPC) to securely distribute aggregation tasks among multiple servers, ensuring privacy despite potential collusion among a subset of them [FMM+21, RSWP23, GMS+23, BMP+24, TXLZ25].
@@
The second (and closely related) challenge is the current lack of global model privacy. Initially, FL focused on enhancing user participation and improving model accuracy rather than safeguarding the model itself. However, with the widespread adoption of FL in various industries, such as the medical sector [RHL+20], the need for privacy protection becomes apparent. Unfortunately, unrestricted access to the aggregated model allows extraction of traces of original training data, necessitating global model privacy [PFA22, BDS+23a, ZSE+24, WCH+24]. So far, only a small number of works have focused on preserving the privacy of the global model in FL [MG19, SPT+21]. Most of these works use homomorphic encryption (HE) schemes [AAUC18], which primarily address semi-honest corruptions. Moreover, several works do not consider the collusion of clients with the aggregator and thereby provide a weaker notion of security. Furthermore, organizations coordinating FL may have a vested interest in keeping the global model private—even from the contributing clients—as revealing the model could pose business risks [DXD+23].
@@
The third challenge arises when corrupted clients intentionally manipulate the locally trained model to reduce the system’s accuracy or introduce backdoors to extract sensitive information in future rounds [WSR+20, XHCL20, ZPS+22]. This method of model poisoning is more potent than data poisoning [SHKR22, XFG25], which involves manipulating the dataset. Although numerous defenses have been proposed against model poisoning [SH21, CFLG21, ZCJG22], the most effective remain impractical for large-scale deployment due to their high computation and communication costs [NRC+22, RSWP23, TXLZ25]. Additionally, recent research has introduced even more potent model poisoning strategies that can circumvent state-of-the-art defenses [SH21, GGP24, XFG25]. For instance, PoisonedFL [XFG25] breaks eight state-of-the-art defenses without any knowledge of the honest clients’ models, highlighting the urgent need for new defense mechanisms. To address corrupted clients, we limit their attack capabilities to data poisoning rather than the more severe model poisoning. This allows us to implement more moderate defenses compared to existing schemes (cf. §5 for details).
@@
Besides these three challenges, a large-scale, worldwide FL deployment needs to account for client heterogeneity in terms of bandwidth, computational power, jurisdiction, and trust assumptions. As we will elaborate in §1.1, trust in government and private entities varies widely, influencing how data is shared and used.
@@
Although various works have addressed the aforementioned issues individually or in combination, a comprehensive solution remains elusive. Therefore, we propose a unified framework called WW-FL, which enables secure and private distributed machine learning at scale. Departing from traditional FL, we introduce the concept of a trust zone, allowing data to be shared securely within a trusted legal jurisdiction and leveraging privacy-enhancing technologies like MPC. This approach supports edge nodes with limited resources and encourages wider participation, enhancing the system’s overall privacy and performance.
1.1 A Regulatory and Trust Perspective
Beyond the critical security, privacy, and technical deployment challenges that exist in FL and have been discussed so far, it is equally important to consider the regulatory landscape and the varying degrees of trust among participating entities. Specifically, WW-FL aims to address the heterogeneity in trust across entities distributed geographically [OEC22, Sta22]. This variation is not only regional but also institutional, with differing levels of trust in, for example, law enforcement versus judicial bodies [OEC22].
When it comes to private companies, some countries have implemented strict laws to ensure the ethical behavior of entities operating within their borders. Non-compliance can result in substantial fines and even imprisonment [Hus23, Hil23]. Recent examples include Meta receiving a 1.2 billion Euro fine from regulators in Ireland [Hus23] and Didi Global facing a 1.19 billion Euro penalty from the Cyberspace Administration of China [Hil23]. Several companies including well-known TalkTalk and Target either lost their customers’ data in major data breaches or were accused of not keeping customer information safe. This led to a lack of trust among their clients and a significant drop in revenue [Eri22, Hib22].
Given this disparity in trust, consider a consortium comprising three companies interested in training a machine learning (ML) model for medical diagnostics. Numerous studies have demonstrated the effectiveness of ML models in diagnosing various diseases, such as lung cancer, tumor detection, and liver segmentation [RHJ+24]. However, these studies emphasize the need for diverse patient data. Simultaneously, the sensitivity and privacy of the training data must be preserved.
To address this, the consortium may collaborate with various federal and local government agencies to obtain relevant training data and potentially compensating data providers. However, since the model is trained on sensitive medical data, the consortium members do not want the trained model to be disclosed to anyone. Additionally, the consortium should have the ability to use the trained model for inference while preserving privacy. This approach not only enhances the system’s privacy by safeguarding against breaches that might affect individual companies but also safeguards the financial investments of these companies by preventing unauthorized use of their trained model by third parties.
At first glance, implementing a PPML solution may appear simplistic. In PPML, data owners securely distribute their data among the three companies within the consortium, ensuring that no individual company can access the data in the clear. The consortium employs MPC techniques to privately train a model on the shared data. However, this approach has several drawbacks. One significant drawback is the substantial computational power required within the consortium to handle the training process, particularly when dealing with millions of data owners in a global deployment. Furthermore, this approach necessitates a high level of trust from all data owners, as they must share their data among the consortium’s member companies. This trust requirement can be a significant hurdle to adoption. Additionally, there may be situations where data owners are willing to share their data but cannot do so due to government regulations preventing data from leaving their jurisdiction. This restriction not only limits data availability but also renders this approach unsuitable for our specific use case.
The next approach we can consider is using a traditional FL with secure aggregation. While this method has the potential to address the scalability problem, it comes with a significant drawback. In this approach, the model during training is distributed to the data owners. Although there are works like [FTP+21, SPT+21, XHX+23], which protect the privacy of the global model using homomorphic encryption techniques, they may not be well-suited for our cross-device setting with a large number of data owners. Additionally, these techniques necessitate interaction among the clients and require extra measures for handling dropouts.
In light of the limitations of current methods, we make an adjustment to the concept of FL. Specifically, we expand the notion of a trust zone, allowing data to be moved from the data owner to this larger, still-trusted area. For instance, a user in our scenario can securely share their data among three entities: the police, the court, and a private company, all operating within the same legal jurisdiction as the data owner. These selected entities will collaboratively train the model using privacy-enhancing technologies like MPC and HE. Furthermore, this approach has the added benefit of supporting edge nodes with limited computational resources, allowing them to contribute to the system. Importantly, since data owners can be assured that their data remains within their designated trust zone, this will likely encourage more users to participate, leading to a larger pool of training data and, ultimately, a better-trained model while guaranteeing global privacy.
Finally, it is important to note that the concept of a trust-zone can differ depending on the specific region or context. For example, in one country, it may be relatively straightforward to identify three entities that data owners trust not to collude with each other. In contrast, there may be situations where it is challenging to find three or four entities that can provide an absolute guarantee of non-collusion. However, in such cases, data owners may be willing to accept a weaker assurance that not all entities will collude. We can accommodate this heterogeneity in trust by performing PPML training independently for each of these scenarios, using the appropriate MPC protocol, and then combining their results through secure aggregation using global servers.
1.2 Our Contributions
@@
In this paper, we initiate a discussion on novel approaches towards secure and private large-scale (cross-device) federated learning to address the challenges outlined above. Towards this goal, we propose a unified framework called “worldwide-FL” (WW-FL). Our framework is based on a novel abstraction that also captures existing hierarchical FL architectures in a hybrid manner as visualized in Fig. 1 and detailed in §3.5.
Figure 1: Visual representation of deploying our secure world-wide FL (WW-FL) framework. In this illustration, clients connect to local MPC clusters, which in turn are connected to three global servers (this can be varied as per requirements) strategically positioned across continents. Fig. 2 provides a detailed illustration of WW-FL, showcasing the roles of involved entities.
@@
Briefly, in WW-FL, clients use secret-sharing to securely outsource training data to distributed cluster servers based on MPC. The clients then can leave and return sporadically to provide more training data – this makes our framework robust against real-world issues like drop-outs, requires no interaction between clients, and resource-constrained (mobile or edge) devices only have very little workload. The models trained by the cluster servers with MPC-based PPML techniques are then aggregated across training clusters using one or multiple levels of distributed aggregators. For secure distributed aggregation, we again utilize MPC. After aggregation, models are not converted to plaintext but are returned to the training clusters for the next training iteration in secret-shared form. When training is completed, known secure inference protocols can be used to allow private queries [NC23, MWCB23] in a controlled (potentially rate-limited) way. Our architecture addresses issues with single aggregators by relying on a distributed aggregator and notably achieves global model privacy due to a novel combination of FL with PPML techniques.
@@
We observe that a unique property of our framework is the strictly limited attack surface: similar to PPML, malicious clients are restricted to weaker data-poisoning attacks as there is no possibility to access and manipulate the model itself. As defending against data poisoning in PPML is highly non-trivial, we show that state-of-the-art data-poisoning attacks in the suggested hierarchical configuration are less effective than in plain FL. Furthermore, we consider different robust aggregation schemes to further mitigate the effect of such attacks. For this, we additionally propose new heuristics that improve the efficiency in MPC. Specifically, we propose a lightweight variant of the “Trimmed Mean” [YCRB18] defense for our setting with only data-poisoning attacks.
@@
Finally, we systematically evaluate possible instantiations of WW-FL in terms of performance (both in terms of achievable accuracy as well as overhead due to MPC) and resilience against poisoning attacks. For this, we implement a prototype based on PyTorch and Meta’s CrypTen framework [KVH+21] for instantiating MPC. We then evaluate the performance when training neural networks for standard image classification tasks in realistic network settings and using GPU-accelerated AWS EC2 instances. This way, we show that the computation and communication overhead induced by MPC is feasible even for large-scale deployments. For example, one round of training LeNet privately with security against semi-honest servers on the MNIST dataset for 5 epochs with batch size 80 takes less than 5 minutes for a cluster with 10 clients, and global aggregation among 10 clusters takes less than 1 second. Moreover, we find that the impact of MPC-induced inaccuracies is less than 0.1%, and models trained with WW-FL converge significantly faster than in plain FL.
[…]
@@
    We propose a new generic (hierarchical) FL framework called WW-FL that achieves global model privacy, supports resource-constrained mobile or edge devices, and significantly limits the attack surface for malicious clients to data poisoning.
@@
  • $\bullet$
    We experimentally compare WW-FL to state-of-the-art collaborative learning frameworks with global model privacy, demonstrating superior scalability and performance.
2 Preliminaries and Related Work
@@
This section provides a brief background on privacy-enhancing technologies like MPC and HE, and an introduction to FL, along with a concise overview of related works. We provide a comparative positioning in Tab. 1 showing that no prior work simultaneously provides global model privacy and protection against malicious clients.
Secure Multi-Party Computation (MPC)
@@
MPC [Yao86, GMW87] enables a set of mutually distrusting parties to evaluate a public function $f()$ on their private data while preserving input data privacy. The corruption among the parties is often modelled via an adversary that may try to take control of the corrupted parties and coordinate their actions. There are various orthogonal aspects of adversarial corruption like honest vs. dishonest majority, semi-honest vs. malicious corruption, etc. [EKR18, Lin20]. MPC is practically efficient when run among a small number of computing parties [PS20, KVH+21, KS22, KPRS22, HSW+25]. In our prototype implementation [^1] Our framework is adaptable to various settings and threat models, as discussed in §3.3. We built our implementation on CrypTen due to its maturity and ML-friendly design. of WW-FL, we utilize Meta’s CrypTen framework [KVH+21], which for efficiency reasons operates in a semi-honest two-party setting with a trusted third “helper” party that generates correlated randomness [RWT+18, CCPS19, PSSY21].
Homomorphic Encryption (HE)
@@
HE schemes [RAD78, Gen09] enable computation on encrypted data without the need for decryption. Additively homomorphic encryption (AHE) is a widely used method that allows for the generation of a new ciphertext representing the sum of multiple plaintexts through operations on their corresponding ciphertexts [Pai99]. In scenarios involving multiple parties, recent multi-party homomorphic encryption (MHE) schemes can reduce the communication complexity of MPC, but have high computation overhead [MTPBH21]. Furthermore, they are often secure only against semi-honest corruptions, whereas [CMS+23] is an exception. A comprehensive survey of various HE schemes is given in [AAUC18].
Differential Privacy (DP)
@@
The concept of DP [DR14] is based on the idea of adding noise to data in order to reduce information leakage when sharing it, while still allowing for meaningful computations to be carried out on the data. Though DP techniques offer some protection against attacks in federated learning, they also reduce the utility of the data and the resulting ML model. Additionally, achieving robust and accurate models necessitates significant privacy budgets, leaving the actual level of privacy achieved in practice uncertain [OA22]. We refer to [DP20] for more details on DP.
Federated Learning (FL)
@@
Unlike conventional PPML techniques, FL [KMY+16, MMR+17] enables the training of an ML model on distributed data by allowing each device to train the model locally using its own data. In each iteration, these local models are sent to a central server (also called an aggregator), where they are aggregated to form a global model. At a high level, an FL scheme iterates through the following steps:
@@
Table 1: Comparison of WW-FL and previous works. Notations: $\mathcal{S}$ – Aggregation Server(s), $\mathcal{C}$ – Client, GM – Global Model, LM – Local Model, AHE/MHE – Additively/Multi-Party Homomorphic Encryption, ZKP – Zero-Knowledge Proof, MPC – Secure Multi-party Computation, DP – Differential Privacy. In terms of privacy, we focus on the protection of the global model on both server and client side as well as the protection of individual client models against a curious or malicious server. Given the extensive literature, this comparison focuses on one representative work per category.
Categories Representative Work(s) Privacy Method Privacy ($\mathcal{S}$) Privacy ($\mathcal{C}$) Defense Cross Device No Client Interaction Dropout Handling
GM LM GM
Aggregation (Plain) [MMR+17]
Aggregation (Robust) [BMGS17]
[CFLG21, ZWP+23]
Secure Aggregation (Single $\mathcal{S}$) [BIK+17] Masking
[MG19] AHE
[GJS+23] Masking+ZKP
[MYP24, LBG+25] Masking
Secure Aggregation (Multi $\mathcal{S}$) [FMM+21] MPC
[…]
[NRC+22] MPC
[RSWP23] MPC $\ddagger$
[TXLZ25] MPC $\ddagger\star$
Hierarchical FL (HFL) [BEG+19] Masking
[WXL+21]
[Yan21] DP
[FHC+24]
[WHZ+25] Masking
[…]
  • $\ast$ Represents collaborative learning in the $N$-party setting, rather than the Federated Learning (FL) setting. SPINDLE [FTP+21] supports only generalized linear models and lacks protocols for training complex models like NNs; POSEIDON [SPT+21] and HERCULES [XHX+23] use activation function approximations, reducing final accuracy.
  • $\ddagger$ Limited to defenses that operate independently on each client’s gradient; excludes aggregation methods like trimmed mean or median [YCRB18].
  • $\star$ Cannot detect malformed gradients that lie within the norm bound (e.g., shrunk gradients).
  • $\dagger$} WW-FL is a generic design that can be implemented with various secure computation protocols, such as MPC or HE.
Secure Aggregation (SA)
@@
Standard FL’s privacy issues [GBDM20, ZSE+24, WBS+24, DHW+25] led to single-server secure aggregation protocols (Single $\mathcal{S}$[BIK+17, BBG+20, MOJC23] that use masking techniques or HE [Rub96, PAH+18, MG19]. However, complete privacy for the global model remains elusive, posing risks if clients and aggregators collude. Moreover, recent research has shown that a single malicious server can reconstruct individual training data points from users’ local models even when using secure aggregation protocols [PFA22, BDS+23a, ZSE+24, WCH+24]. Such vulnerabilities against colluding parties are addressed by multi-server SA (Multi $\mathcal{S}$[FMM+21, DCL+21, NRC+22, RSWP23, TXLZ25] and techniques like multi-party homomorphic encryption (MHE) [FTP+21, SPT+21, XHX+23], albeit with scalability challenges. Works like [DCL+21] and [NRC+22] combine HE and MPC techniques to achieve private and robust FL, albeit at the cost of considerable computation and communication overhead in cross-device scenarios comprising hundreds to thousands of clients [SHKR22]. In WW-FL, we employ a multi-server aggregation scheme, and mitigate scalability issues of existing works, especially for a heterogeneous cross-device setting.
Collaborative Learning with Global Model Privacy
Several prior works focus explicitly on preserving the privacy of the global model during collaborative learning. These approaches primarily rely on HE, particularly multi-party homomorphic encryption (MHE) [MTPBH21], to securely aggregate model updates while maintaining privacy across multiple clients. MHE enables distributed and fully decentralized training through direct interactions among data owners in settings where non-colluding servers are unavailable. Spindle [FTP+21] employs MHE to support secure distributed training, specifically for logistic regression models. However, Spindle is limited to linear models, which restricts its applicability to more complex architectures. Poseidon [SPT+21] extends MHE to 3-layer neural networks by leveraging multi-party CKKS [MTPBH21]. In this framework, multiple entities collaboratively generate a shared secret key, enabling encrypted computations under a common public key. To maintain computational efficiency, Poseidon replaces non-polynomial functions, such as the commonly used ReLU activation function, with polynomial approximations. However, these polynomial approximations introduce accuracy trade-offs that may significantly impact model performance. To address these computational inefficiencies, Hercules [XHX+23] was proposed as an enhancement of Poseidon. Hercules introduces optimized parallel homomorphic operations, including SIMD-enabled matrix computations and improved polynomial approximations for RELU. While Hercules offers higher computational performance than Poseidon, it continues to suffer from accuracy degradation due to the use of polynomial approximations for non-linear layers.
Despite these advancements, all three MHE-based frameworks – Spindle [FTP+21], Poseidon [SPT+21], and Hercules [XHX+23] – exhibit notable limitations. The reliance on polynomial approximations for non-linear functions, particularly ReLU, inherently reduces model accuracy [RRK+20]. Moreover, these frameworks lack dedicated robustness mechanisms against poisoning attacks, without which a single malicious participant can substantially degrade the global model accuracy [FCJG20]. Scalability also remains a critical limitation: the experiments in these works typically involve no more than 50 parties, raising concerns about their feasibility in larger, real-world collaborative scenarios. Additionally, these approaches require interactive key setup among clients for encrypting each model and cooperative decryption of the final aggregated model, which introduces further complexity and communication overhead.
[…]
Hierarchical Federated Learning (HFL)